40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1558     Entry time: Sat Aug 1 09:24:58 2015
Author: Aidan, Antonio 
Type: Summary 
Category: BEAT 
Subject: Lasers locked to cavities - no beat - polarization? Transmission issues 

Both lasers have been locked to the cavities for 24 hours. The slow control of the frequency is handed off to the PID loop. Antonio and I observed strange behaviour on the DC value of the cavity transmission.

As Evan had noted before, there are two polarizations that will resonate and they're about 3MHz apart (if I remember correctly). We can see these on the DC photodiodes on transmission (the ISS PD and the RF DC output). One peak is large and the other much smaller. However, when we have large transmission onto the RF photodiode we have small transmission onto the ISS PD and vice versa. It's likely we have a pick-off optic with strong polarization selectivity.

We couldn't find the beat yesterday or Thursday.

ELOG V3.1.3-